OpenStudy (anonymous):

The demand equation for the Drake GPS Navigator is x + 4p – 800 = 0, where x is the quantity demanded per week and p is the wholesale unit price in dollars. The supply equation is x – 20p +1000 = 0, where x is the quantity the supplier will make available in the market when the wholesale price is p dollars each. Find the equilibrium quantity for the GPS Navigators. need the answer please for a test

OpenStudy (anonymous):

demand: x + 4p – 800 = 0 x = 800 -4p and supply: x – 20p +1000 = 0 x= 20p -1000

OpenStudy (anonymous):

we need equilibrium, which means supply = demand 800 - 4p = 20p -1000

OpenStudy (anonymous):

24p = 1800

OpenStudy (anonymous):

p = 1800/24 = 75

OpenStudy (anonymous):

75 is the answer?

OpenStudy (anonymous):

do you understand the method?

OpenStudy (anonymous):

75 is wrong

OpenStudy (anonymous):

75 is not the answer

OpenStudy (anonymous):

p = 75, we now use p to find x

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

x= 20p -1000 x= 20*75 - 1000 = 500

OpenStudy (anonymous):

we could have found x straight away but i thought the algebra was easier this way

OpenStudy (anonymous):

x=500=500?

OpenStudy (anonymous):

yes. if i have understood the question properly

OpenStudy (anonymous):

answer is 500?

OpenStudy (anonymous):

yes